# Adaptive Off-Policy Inference for M-Estimators Under Model Misspecification

### v1

@@ B.3 Proof of Proposition 1
$\mathbb{E}\left[s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1},X_{t}\right]$ $=\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{t})\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\mid X_{t}\right]$
$=\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]:=\bar{\mu}(X_{t}).$
On the other hand, the conditional variance given $X_{t}$ can be written as:
$\mathrm{Var}\left(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1},X_{t}\right)$ $=\mathrm{Var}\Bigg(\sum_{a\in\mathcal{A}}\pi_{e}(a\mid X_{t})\frac{\mathbbm{1}_{A_{t}=a}}{\pi_{t}(a\mid X_{t})}\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})\,\Big|\,\mathcal{H}_{t-1},X_{t}\Bigg)$
$=\sum_{a\in\mathcal{A}}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\underbrace{\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{2}\mid X_{t}]}_{v^{\star}(X_{t},a)+\mu^{\star}(X_{t},a)^{2}}-\underbrace{\Big(\sum_{a\in\mathcal{A}}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)^{2}}_{\text{mean correction}}.$
Here we used that $\dot{m}_{\theta}(a,X_{t},f_{t}(a,X_{t}))\in\sigma(X_{t},\mathcal{H}_{t-1})$.
@@
$\mathrm{Var}(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1})$ $=\mathrm{Var}\Big(\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mid\mathcal{H}_{t-1}\Big)$
$\quad+\mathbb{E}\Bigg[\sum_{a\in\mathcal{A}}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\left(v^{\star}(X_{t},a)+\mu^{\star}(X_{t},a)^{2}\right)-\Big(\sum_{a\in\mathcal{A}}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)^{2}\;\Bigg|\;\mathcal{H}_{t-1}\Bigg]$
$=\mathrm{Var}\Big[\mathbb{E}_{A_{t}\sim\pi_{e}}[\dot{m}_{\theta^{\star}}(A_{t},X_{t},Y_{t})\mid X_{t}]\mid\mathcal{H}_{t-1}\Big]+\mathbb{E}\Big[\mathrm{Var}_{A\sim\tilde{\pi}_{e}}(\dot{m}_{\theta^{\star}}(A,X_{t},Y_{t}(a))\mid X_{t})\,\Big|\,\mathcal{H}_{t-1}\Big],$
[…]
@@ B.4 Proof of Proposition 2
$\sum_{a\in\mathcal{A}}\pi_{e}(a|X_{t})\hat{\nu}_{t}(a,X_{i})-\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})\right]$ $=\sum_{a\in\mathcal{A}}\pi_{e}(a|X_{t})\hat{\nu}_{t}(a,X_{i})-\sum_{a\in\mathcal{A}}\pi_{e}(a|X_{t})\mathbb{E}\left[\dot{m}_{\theta}(X_{t},a,Y_{t})|X_{t}\right]$
$=\sum_{a\in\mathcal{A}}\pi_{e}(a|X_{t})\left(\hat{\nu}_{t}(a,X_{i})-\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|A_{t}=a,X_{t}\right]\right)$
$\leq K\left(\hat{\nu}_{t}(a,X_{i})-\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|A_{t}=a,X_{t}\right]\right)$
@@
For ease of notation, let us write $\nu^{\star}_{t}(X_{t}):=\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right]$.
Let us consider the first term in the expression $\frac{1}{n}\left(\sum_{X_{i}\in\tilde{X}}\hat{\nu}_{t}(X_{i})^{2}-\bar{\nu}\right)\left(\sum_{X_{i}\in\tilde{X}}\hat{\nu}_{t}(X_{i})^{2}-\bar{\nu}\right)^{T}$. Let us substitute $\hat{\nu}_{t}(X_{i})=\hat{\nu}_{t}(X_{i})-\nu^{\star}_{t}(X_{t})+\nu^{\star}_{t}(X_{t})$. We then obtain the result that the above expression is equal to
@@
Now, let us consider the term
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}\frac{\pi_{e}(A_{t}=a|X_{t})^{2}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}h_{t}(X_{i},a)-\mathbb{E}\left[\sum_{a\in\mathcal{A}}\frac{\pi_{e}(A_{t}=a|X_{t})^{2}}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]|\mathcal{H}_{t-1}\right].$
@@
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\pi_{e}(A_{t}=a|X_{t})^{2}$ $\sum_{a\in\mathcal{A}}\frac{\left(h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]\right)}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}$
$\leq\frac{1}{nC_{1}}\left[\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}h_{t}(X_{i},a)\right]-\frac{1}{C_{1}}\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right].$
To show that this term is $o_{p}(1)$, it is sufficient to show that
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]=o_{p}(1)$
for all $a\in\mathcal{A}$. Note that
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)$ $=\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)+\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]$
$\quad-\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]$
@@
Therefore, $\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]=o_{p}(1)$ immediately by the law of large numbers.
Finally, consider the term
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}g_{t}(X_{i},a)-\mathbb{E}\left[\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}g_{t}(X_{i},a)\mid\mathcal{H}_{t-1}\right].$
As before, write
$g_{t}(X_{i},a)=g_{t}(X_{i},a)-\mathbb{E}[g_{t}(X_{i},a)\mid X_{i}]+\mathbb{E}[g_{t}(X_{i},a)\mid X_{i}].$
Then the term becomes
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}\big(g_{t}(X_{i},a)-\mathbb{E}[g_{t}(X_{i},a)\mid X_{i}]\big)$
$\quad+\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}\mathbb{E}[g_{t}(X_{i},a)\mid X_{i}].$
The first sum is $o_{p}(1)$, and the second sum converges to its expectation by the weak law of large numbers. Hence,
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}g_{t}(X_{i},a)-\mathbb{E}\left[\sum_{a\in\mathcal{A}}\pi_{e}(A_{t}=a\mid X_{i})^{2}g_{t}(X_{i},a)\mid\mathcal{H}_{t-1}\right]=o_{p}(1).$

### v2

@@ B.3 Proof of Proposition 1
For ease of notation, we denote the quantities $\pi_{t}(a\mid X_{t}):=\mathbb{P}(A_{t}\mid X_{t},\mathcal{H}_{t-1})$, $\mu^{\star}(X_{t},a):=\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\mid X_{t}]$, and $v^{\star}(X_{t},a):=\text{Var}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\mid X_{t}]$.
@@
$\mathbb{E}\left[s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1},X_{t}\right]$ $=\mathbb{E}\Bigg[\sum_{a=1}^{K}\pi_{e}(A_{t}=a\mid X_{t})\Big(\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a))$
$\qquad\qquad+\mathbbm{1}_{A_{t}=a}\frac{\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})-\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a)}{\pi_{t}(a\mid X_{t})}\Big)\,\Big|\,\mathcal{H}_{t-1},X_{t}\Bigg]$
$=\sum_{a=1}^{K}\pi_{e}(A_{t}=a\mid X_{t})\mathbb{E}\left[\frac{\mathbbm{1}_{A_{t}=a}}{\pi_{t}(a\mid X_{t})}\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\mid X_{t},\mathcal{H}_{t-1}\right]$
$\qquad\qquad+\sum_{a=1}^{K}\pi_{e}(A_{t}=a\mid X_{t})\mathbb{E}\left[\left(1-\frac{\mathbbm{1}_{A_{t}=a}}{\pi_{t}(a\mid X_{t})}\right)\dot{m}_{\theta^{\star}}(X_{t},a,f_{t}(X_{t},a)\mid X_{t},\mathcal{H}_{t-1}\right]$
$=\sum_{a=1}^{K}\pi_{e}(A_{t}=a\mid X_{t})\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\mid X_{t}\right]$
$=\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right].$
On the other hand, note that $\dot{m}_{\theta}(a,X_{t},f_{t}(a,X_{t}))\in\sigma(X_{t},\mathcal{H}_{t-1})$. Therefore, we can rewrite the conditional variance given $X_{t}$ as:
$\mathrm{Var}\left(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1},X_{t}\right)$ $=\mathrm{Var}\Bigg(\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\frac{\mathbbm{1}_{A_{t}=a}}{\pi_{t}(a\mid X_{t})}\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t})\,\Big|\,\mathcal{H}_{t-1},X_{t}\Bigg)$
$=\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{T}\mid X_{t}]$
$\qquad-\Big(\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)\Big(\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)^{T}$
$=\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{T}\mid X_{t}]$
$\qquad-\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}.$
@@
$\mathrm{Var}(s_{t,\theta^{\star}}\mid\mathcal{H}_{t-1})$ $=\mathrm{Var}\Big(\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\Big)$
$\qquad+\mathbb{E}\Bigg[\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{T}\mid X_{t}]\mid\mathcal{H}_{t-1}\Bigg]$
$\qquad-\mathbb{E}\left[\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}\right].$
@@ B.4 Proof of Proposition 2
For ease of notation, let us write $\nu^{\star}_{t}(X_{t}):=\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right]$. We will tackle this proof in three parts:
  1. 1.
    Showing $\frac{1}{n-1}\Biggl(\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})-\bar{\nu}\Biggr)\Biggl(\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})-\bar{\nu}\Biggr)^{T}-\mathrm{Var}\left(\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right]\right)=o_{p}(1)$;
  2. 2.
    Showing
    $\frac{1}{n}\sum_{X_{i}\in\mathcal{\tilde{X}}}\sum_{a=1}^{K}\frac{\pi_{e}(A_{t}=a\mid X_{i})^{2}\,h_{t}(X_{i},a)}{\mathbb{P}(A_{t}=a\mid X_{i},\mathcal{H}_{t-1})}$
    $\qquad-\mathbb{E}\Bigg[\sum_{a=1}^{K}\frac{\pi_{e}(a\mid X_{t})^{2}}{\pi_{t}(a|X_{t})}\mathbb{E}[\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))\dot{m}_{\theta^{\star}}(X_{t},a,Y_{t}(a))^{T}\mid X_{t}]\mid\mathcal{H}_{t-1}\Bigg]=o_{p}(1);$
  3. 3.
    Showing
    $\frac{1}{n}\sum_{X_{i}\in\mathcal{\tilde{X}}}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}-\mathbb{E}\Bigg[\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}\mid\mathcal{H}_{t-1}\Bigg]=o_{p}(1).$
Adding term (1)-(3) together will demonstrate that $\left\lVert\hat{V}_{t}-V_{t,\theta^{\star}}\right\rVert_{\text{op}}\overset{p}{\to}0$.
Showing (1)
@@
$\hat{\nu}_{t}(X_{i})-\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|X_{t}\right]$ $=\sum_{a=1}^{K}\pi_{e}(a|X_{t})g_{t}(a,X_{i})-\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})\right]$
$=\sum_{a=1}^{K}\pi_{e}(a|X_{t})g_{t}(a,X_{i})-\sum_{a=1}^{K}\pi_{e}(a|X_{t})\mathbb{E}\left[\dot{m}_{\theta}(X_{t},a,Y_{t})|X_{t}\right]$
$=\sum_{a=1}^{K}\pi_{e}(a|X_{t})\left(g_{t}(a,X_{i})-\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|A_{t}=a,X_{t}\right]\right)$
$\leq K\left(g_{t}(a,X_{i})-\mathbb{E}\left[\dot{m}_{\theta}(X_{t},A_{t},Y_{t})|A_{t}=a,X_{t}\right]\right)$
@@
The final line holds by 9.
Let us consider the first term in the expression $\frac{1}{n}\left(\sum_{X_{i}\in\tilde{X}}\hat{\nu}_{t}(X_{i})-\bar{\nu}\right)\left(\sum_{X_{i}\in\tilde{X}}\hat{\nu}_{t}(X_{i})^{2}-\bar{\nu}\right)^{T}$. Let us substitute $\hat{\nu}_{t}(X_{i})=\hat{\nu}_{t}(X_{i})-\nu^{\star}_{t}(X_{t})+\nu^{\star}_{t}(X_{t})$. We then obtain the result that the above expression is equal to
@@
Showing (2)
Now, let us consider the expression
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}\sum_{a=1}^{K}\frac{\pi_{e}(A_{t}=a\mid X_{t})^{2}}{\mathbb{P}\!\left(A_{t}=a\mid X_{t},\mathcal{H}_{t-1}\right)}\,h_{t}(X_{i},a)$
$\quad-\mathbb{E}\Bigg[\sum_{a=1}^{K}\frac{\pi_{e}(A_{t}=a\mid X_{t})^{2}}{\mathbb{P}\!\left(A_{t}=a\mid X_{t},\mathcal{H}_{t-1}\right)}\,\mathbb{E}\Big[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}\;\Big|\;X_{t},A_{t}=a\Big]\;\Big|\;\mathcal{H}_{t-1}\Bigg].$
@@
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}$ $\sum_{a=1}^{K}\pi_{e}(A_{t}=a|X_{t})^{2}\frac{\left(h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]\right)}{\mathbb{P}\left(A_{t}=a|X_{t},\mathcal{H}_{t-1}\right)}$
$\leq\frac{C_{1}}{n}\left[\sum_{X_{i}\in\tilde{X}}\sum_{a=1}^{K}h_{t}(X_{i},a)\right]-C_{1}\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right].$
To show that this expression is $o_{p}(1)$, it is sufficient to show that
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]=o_{p}(1).$
for all $a\in\mathcal{A}$. Note that by Assumption 9,
$\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)$ $=\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)+\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]$-
$\qquad\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]$
@@
Therefore, $\frac{1}{n}\sum_{X_{i}\in\tilde{X}}h_{t}(X_{i},a)-\mathbb{E}\left[\mathbb{E}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})^{T}|X_{t},A_{t}=a\right]\right]=o_{p}(1)$ by the weak law of large numbers.
Showing (3)
Finally, consider the expression
$\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}-\Big(\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)\Big(\sum_{a=1}^{K}\pi_{e}(a\mid X_{t})\mu^{\star}(X_{t},a)\Big)^{T}.$ (10)
Let $\Delta_{t}(x):=\widehat{\nu}_{t}(x)-\nu_{t}^{\star}(x)$ which we already showed was $o_{p}(1)$. Now, we can write:
$\widehat{\nu}_{t}(X_{i})\widehat{\nu}_{t}(X_{i})^{\!\top}$ $=\big(\Delta_{t}(X_{i})+\nu_{t}^{\star}(X_{i})\big)\big(\Delta_{t}(X_{i})+\nu_{t}^{\star}(X_{i})\big)^{\!\top}$
$=\Delta_{t}(X_{i})\Delta_{t}(X_{i})^{\!\top}\;+\;\Delta_{t}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}\;+\;\nu_{t}^{\star}(X_{i})\Delta_{t}(X_{i})^{\!\top}\;+\;\nu_{t}^{\star}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}.$
Subtracting the last term on both sides yields the identity
$\widehat{\nu}_{t}(X_{i})\widehat{\nu}_{t}(X_{i})^{\!\top}-\nu_{t}^{\star}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}=\Delta_{t}(X_{i})\Delta_{t}(X_{i})^{\!\top}+\Delta_{t}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}+\nu_{t}^{\star}(X_{i})\Delta_{t}(X_{i})^{\!\top}.$
Each individual term is $o_{p}(1)$ because $\nu_{t}^{\star}(X_{i})$ is integrable. Therefore, we have that Equation 10 is $o_{p}(1)$. Now, consider the expression.
$\frac{1}{n}\sum_{i=1}^{n}\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}-\mathbb{E}\left[\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}\mid\mathcal{H}_{t-1}\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{\nu}_{t}(X_{i})\hat{\nu}_{t}(X_{i})^{T}-\nu_{t}^{\star}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}\right)$
$\qquad+\frac{1}{n}\sum_{i=1}^{n}\nu_{t}^{\star}(X_{i})\nu_{t}^{\star}(X_{i})^{\!\top}-\mathbb{E}\left[\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]\mathbb{E}_{A_{t}\sim\pi_{e}}\left[\dot{m}_{\theta^{\star}}(X_{t},A_{t},Y_{t})\mid X_{t}\right]^{T}\mid\mathcal{H}_{t-1}\right].$
The first term is $o_{p}(1)$ by the arguments given above, the second term is $o_{p}(1)$ by the weak law of large numbers.